OpenStudy (anonymous):

What is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9,374? 11 11.5 12 12.5

OpenStudy (campbell_st):

use \[9374 =\frac{43}{2}[2\times( -13) + (43-1)4d]\] simplify the equation \[ 9374 = -559 +903 d\] now solve for d

OpenStudy (anonymous):

11?

OpenStudy (campbell_st):

thats what I calculated

OpenStudy (anonymous):

thanks!